\begin{thebibliography}{11}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[ai et~al.(2025)ai, Teng, Jia, Sun, Li, Li, Tang, Han, Zhang, Zhang, Luo, Kang, Sun, Cao, Huang, Lin, Fang, Tao, Zhang, Wang, Liu, Shi, Su, Sun, Pan, Wang, Sheng, Cui, Hu, Yan, Yin, Zhang, Liu, Yin, Yang, Song, Hu, Zhang, and Li]{magi1}
Sand. ai, Hansi Teng, Hongyu Jia, Lei Sun, Lingzhi Li, Maolin Li, Mingqiu Tang, Shuai Han, Tianning Zhang, W.~Q. Zhang, Weifeng Luo, Xiaoyang Kang, Yuchen Sun, Yue Cao, Yunpeng Huang, Yutong Lin, Yuxin Fang, Zewei Tao, Zheng Zhang, Zhongshu Wang, Zixun Liu, Dai Shi, Guoli Su, Hanwen Sun, Hong Pan, Jie Wang, Jiexin Sheng, Min Cui, Min Hu, Ming Yan, Shucheng Yin, Siran Zhang, Tingting Liu, Xianping Yin, Xiaoyu Yang, Xin Song, Xuan Hu, Yankai Zhang, and Yuqiao Li.
\newblock Magi-1: Autoregressive video generation at scale, 2025.

\bibitem[Assran et~al.(2025)Assran, Bardes, Fan, Garrido, Howes, Mojtaba, Komeili, Muckley, Rizvi, Roberts, Sinha, Zholus, Arnaud, Gejji, Martin, Hogan, Dugas, Bojanowski, Khalidov, Labatut, Massa, Szafraniec, Krishnakumar, Li, Ma, Chandar, Meier, LeCun, Rabbat, and Ballas]{vjepa2}
Mido Assran, Adrien Bardes, David Fan, Quentin Garrido, Russell Howes, Mojtaba, Komeili, Matthew Muckley, Ammar Rizvi, Claire Roberts, Koustuv Sinha, Artem Zholus, Sergio Arnaud, Abha Gejji, Ada Martin, Francois~Robert Hogan, Daniel Dugas, Piotr Bojanowski, Vasil Khalidov, Patrick Labatut, Francisco Massa, Marc Szafraniec, Kapil Krishnakumar, Yong Li, Xiaodong Ma, Sarath Chandar, Franziska Meier, Yann LeCun, Michael Rabbat, and Nicolas Ballas.
\newblock V-jepa 2: Self-supervised video models enable understanding, prediction and planning, 2025.

\bibitem[Bar-Tal et~al.(2024)Bar-Tal, Chefer, Tov, Herrmann, Paiss, Zada, Ephrat, Hur, Liu, Raj, Li, Rubinstein, Michaeli, Wang, Sun, Dekel, and Mosseri]{lumiere}
Omer Bar-Tal, Hila Chefer, Omer Tov, Charles Herrmann, Roni Paiss, Shiran Zada, Ariel Ephrat, Junhwa Hur, Guanghui Liu, Amit Raj, Yuanzhen Li, Michael Rubinstein, Tomer Michaeli, Oliver Wang, Deqing Sun, Tali Dekel, and Inbar Mosseri.
\newblock Lumiere: A space-time diffusion model for video generation, 2024.

\bibitem[Blattmann et~al.(2023)Blattmann, Dockhorn, Kulal, Mendelevitch, Kilian, Lorenz, Levi, English, Voleti, Letts, Jampani, and Rombach]{svd}
Andreas Blattmann, Tim Dockhorn, Sumith Kulal, Daniel Mendelevitch, Maciej Kilian, Dominik Lorenz, Yam Levi, Zion English, Vikram Voleti, Adam Letts, Varun Jampani, and Robin Rombach.
\newblock Stable video diffusion: Scaling latent video diffusion models to large datasets, 2023.

\bibitem[Brooks et~al.(2024)Brooks, Peebles, Holmes, DePue, Guo, Jing, Schnurr, Taylor, Luhman, Luhman, Ng, Wang, and Ramesh]{sora}
Tim Brooks, Bill Peebles, Connor Holmes, Will DePue, Yufei Guo, Li Jing, David Schnurr, Joe Taylor, Troy Luhman, Eric Luhman, Clarence Ng, Ricky Wang, and Aditya Ramesh.
\newblock Video generation models as world simulators.
\newblock 2024.

\bibitem[Kondratyuk et~al.(2024)Kondratyuk, Yu, Gu, Lezama, Huang, Schindler, Hornung, Birodkar, Yan, Chiu, Somandepalli, Akbari, Alon, Cheng, Dillon, Gupta, Hahn, Hauth, Hendon, Martinez, Minnen, Sirotenko, Sohn, Yang, Adam, Yang, Essa, Wang, Ross, Seybold, and Jiang]{videopoet}
Dan Kondratyuk, Lijun Yu, Xiuye Gu, José Lezama, Jonathan Huang, Grant Schindler, Rachel Hornung, Vighnesh Birodkar, Jimmy Yan, Ming-Chang Chiu, Krishna Somandepalli, Hassan Akbari, Yair Alon, Yong Cheng, Josh Dillon, Agrim Gupta, Meera Hahn, Anja Hauth, David Hendon, Alonso Martinez, David Minnen, Mikhail Sirotenko, Kihyuk Sohn, Xuan Yang, Hartwig Adam, Ming-Hsuan Yang, Irfan Essa, Huisheng Wang, David~A. Ross, Bryan Seybold, and Lu Jiang.
\newblock Videopoet: A large language model for zero-shot video generation, 2024.

\bibitem[Motamed et~al.(2025)Motamed, Culp, Swersky, Jaini, and Geirhos]{physicsIQ}
Saman Motamed, Laura Culp, Kevin Swersky, Priyank Jaini, and Robert Geirhos.
\newblock Do generative video models understand physical principles?, 2025.

\bibitem[{Pika Labs}(2023)]{pika_1.0_2023}
{Pika Labs}.
\newblock Pika 1.0, 2023.

\bibitem[{Runway}(2024)]{runway_gen3_alpha_2024}
{Runway}.
\newblock {Gen-3 Alpha}, 2024.

\bibitem[Wan et~al.(2025)Wan, Wang, Ai, Wen, Mao, Xie, Chen, Yu, Zhao, Yang, Zeng, Wang, Zhang, Zhou, Wang, Chen, Zhu, Zhao, Yan, Huang, Feng, Zhang, Li, Wu, Chu, Feng, Zhang, Sun, Fang, Wang, Gui, Weng, Shen, Lin, Wang, Wang, Zhou, Wang, Shen, Yu, Shi, Huang, Xu, Kou, Lv, Li, Liu, Wang, Zhang, Huang, Li, Wu, Liu, Pan, Zheng, Hong, Shi, Feng, Jiang, Han, Wu, and Liu]{wan2025}
Team Wan, Ang Wang, Baole Ai, Bin Wen, Chaojie Mao, Chen-Wei Xie, Di Chen, Feiwu Yu, Haiming Zhao, Jianxiao Yang, Jianyuan Zeng, Jiayu Wang, Jingfeng Zhang, Jingren Zhou, Jinkai Wang, Jixuan Chen, Kai Zhu, Kang Zhao, Keyu Yan, Lianghua Huang, Mengyang Feng, Ningyi Zhang, Pandeng Li, Pingyu Wu, Ruihang Chu, Ruili Feng, Shiwei Zhang, Siyang Sun, Tao Fang, Tianxing Wang, Tianyi Gui, Tingyu Weng, Tong Shen, Wei Lin, Wei Wang, Wei Wang, Wenmeng Zhou, Wente Wang, Wenting Shen, Wenyuan Yu, Xianzhong Shi, Xiaoming Huang, Xin Xu, Yan Kou, Yangyu Lv, Yifei Li, Yijing Liu, Yiming Wang, Yingya Zhang, Yitong Huang, Yong Li, You Wu, Yu Liu, Yulin Pan, Yun Zheng, Yuntao Hong, Yupeng Shi, Yutong Feng, Zeyinzi Jiang, Zhen Han, Zhi-Fan Wu, and Ziyu Liu.
\newblock Wan: Open and advanced large-scale video generative models.
\newblock \emph{arXiv preprint arXiv:2503.20314}, 2025.

\bibitem[Yang et~al.(2025)Yang, Teng, Zheng, Ding, Huang, Xu, Yang, Hong, Zhang, Feng, Yin, Zhang, Wang, Cheng, Xu, Gu, Dong, and Tang]{cogvideox}
Zhuoyi Yang, Jiayan Teng, Wendi Zheng, Ming Ding, Shiyu Huang, Jiazheng Xu, Yuanming Yang, Wenyi Hong, Xiaohan Zhang, Guanyu Feng, Da Yin, Yuxuan Zhang, Weihan Wang, Yean Cheng, Bin Xu, Xiaotao Gu, Yuxiao Dong, and Jie Tang.
\newblock Cogvideox: Text-to-video diffusion models with an expert transformer, 2025.

\end{thebibliography}

\documentclass[10pt,twocolumn,letterpaper]{article}

\usepackage[pagenumbers]{iccv} 
\usepackage{tabularx}

\definecolor{iccvblue}{rgb}{0.21,0.49,0.74}
\usepackage[pagebackref,breaklinks,colorlinks,allcolors=iccvblue]{hyperref}

\def\paperID{*****} 
\def\confName{ICCV}
\def\confYear{2025}

\title{Improving the physics of video generation with VJEPA-2 reward signal
}

\author{
Jianhao Yuan$^{1,2,}$\thanks{Work done during internship at Meta.} \, Xiaofeng Zhang$^{1,3,4}$ \, Felix Friedrich$^{1}$ \, Nicolas Beltran-Velez$^{1,5,\ast}$ \\ Melissa Hall$^{1}$ \, Reyhane Askari-Hemmat$^{1}$ \, Xiaochuang Han$^{1}$, Nicolas Ballas$^{1}$ \\ Michal Drozdzal$^{1,\dagger}$ \, Adriana Romero-Soriano$^{1,3,6,7,}$\thanks{Senior co-leads.}\\ \\
$^{1}$FAIR, Meta Superintelligence Labs \, $^{2}$University of Oxford\, $^{3}$Mila - Qu\'{e}bec AI Institute \\ $^{4}$Universit\'{e} de Montr\'{e}al \, $^{5}$Columbia University \, $^{6}$McGill University \, $^{7}$Canada CIFAR AI Chair\\
{\tt\small \{adrianars, mdrozdzal\}@meta.com, jianhaoyuan@robots.ox.ac.uk}}

\begin{document}
\maketitle

\emph{This is a short technical report describing the winning entry of the PhysicsIQ Challenge, presented at the Perception Test Workshop at ICCV 2025.}

\section{Introduction}
\label{sec:intro}
State-of-the-art video generative models~\citep{magi1, cogvideox, sora, videopoet, lumiere, wan2025} exhibit severely limited physical understanding, and often produce implausible videos. The Physics IQ benchmark~\citep{physicsIQ} has shown that visual realism does not imply physics understanding. Yet, intuitive physics understanding has shown to emerge from SSL pretraining on natural videos~\citep{vjepa2}. In this report, we investigate whether we can leverage SSL-based video world models to improve the physics plausibility of video generative models. 

In particular, we build ontop of the state-of-the-art video generative model MAGI-1~\citep{magi1} and couple it with the recently introduced Video Joint Embedding Predictive Architecture 2 (VJEPA-2) to guide the generation process~\citep{magi1}. We show that by leveraging VJEPA-2 as reward signal, we can improve the physics plausibility of state-of-the-art video generative models by $\sim6\%$.

\section{Background}
\label{sec:back}

\paragraph{MAGI-1~\cite{magi1}.} MAGI-1 is an autoregressive video generative model that predicts a sequence of video chunks. The model is trained to iteratively denoise each chunk and enables casual temporal video modeling. The model is able to generate videos when conditioned on text, text and image, as well as text and video.

At the inference time, the model uses the following score function to guide the diffusion process: 
\begin{align*}
\nabla_{x_t} \textrm{log} \ p_s (x_t | x_{<t}, \textrm{txt}) = (1- \omega_{<t}) \nabla_{x_t} \textrm{log} \ p(x_t) \\ + \  (\omega_{<t} - \omega_{\textrm{txt}}) \nabla_{x_t} \textrm{log} \ p(x_t | x_{<t}) \\ + \ \omega_{\textrm{txt}} \nabla_{x_t} \textrm{log} \ p(x_t | x_{<t}, \textrm{txt}),   
\label{eq:score}
\end{align*}
 where $x_t$ is the current video chunk, $x_{<t}$ are the past video chunks, txt is the text conditioning, $\omega_{<t}$ is the context guidance weight and $\omega_{\textrm{txt}}$ is the text guidance weight. 

\paragraph{VJEPA-2~\cite{vjepa2}.}  VJEPA-2 is a self-supervised video model which consists of a video encoder, $E_\theta$, and a predictor, $P_{\phi}$, trained with a Internet-scale data using a mask-denoising objective in the representation space.\looseness-1

\section{Method}
\label{sec:method}

We use VJEPA-2 as reward signal during the generation process of the state-of-the-art image-to-video (I2V) and video-to-video (V2V) MAGI-1 model. To do so, we leverage the \emph{surprise score} introduced in~\citep{vjepa2} and compare the generations of MAGI-1 with the VJEPA-2 predictions as follows:\looseness-1

\begin{equation}
S(x_t | x_{<t}, \textrm{txt}) = \textrm{sim}\big(P_{\phi}(E_\theta(x_{<t})), E_\theta(G_\gamma(x_{<t}, \textrm{txt})\big),
\label{eq:surprise}
\end{equation}

where $E_\theta$ and $P_{\phi}$ are the VJEPA-2 encoder and predictor, respectively; $G_{\gamma}$ is the generative model; and $\textrm{sim}$ is the cosine similarity.

We start by altering the score function of MAGI-1 to include signals from VJEPA-2 as follows:

\begin{equation}\label{eq:score}
    \begin{aligned}
\nabla_{x_t} \textrm{log} \ p_s (x_t | x_{<t}, \textrm{txt}) = (1- \omega_{<t}) \nabla_{x_t} \textrm{log} \ p(x_t) \\ + \  (\omega_{<t} - \omega_{\textrm{txt}}) \nabla_{x_t} \textrm{log} \ p(x_t | x_{<t}) \\ + \ \omega_{\textrm{txt}} \nabla_{x_t} \textrm{log} \ p(x_t | x_{<t}, \textrm{txt}) \\ - \ \omega_s \nabla_{x_t} S(x_t | x_{<t},\textrm{txt}) 
\end{aligned}
\end{equation}

Following Eq.~\eqref{eq:score}, we collect 16 video samples per condition. Then, we apply Best of N (BoN) and select the video sample with the lowest average surprise score Eq.~\eqref{eq:surprise}.\looseness-1

\section{Results}
\label{sec:results}

We evaluate our approach on the PhysicsIQ dataset and report the results in Table~\ref{tab:physicsiq}. By leveraging both VJEPA-2's surprise score as guidance and VJEPA-2 signal to select BoN, our method boosts vanilla sampling of MAGI-1 model and reaches a new state-of-the-art result of \textbf{62.64} and \textbf{36.86} on V2V and I2V, respectively.


\begin{table}[ht]
\centering
\caption{Evaluation on \textbf{PhysicsIQ}}
\begin{tabularx}{\linewidth}{lc}
\toprule
& PhysicsIQ Score ($\uparrow$)\\
\midrule
\multicolumn{2}{l}{\emph{V2V Generation}} \\
\midrule
Lumiere~\citep{lumiere}  & 23.00 \\
VideoPoet~\citep{videopoet} & 29.50 \\
MAGI-1~\citep{magi1}   & 56.31 \\
\midrule
MAGI-1 + VJEPA-2 (ours) & \textbf{62.64}	(+6.33) \\
\midrule
\multicolumn{2}{l}{\emph{I2V Generation}} \\
\midrule
Sora~\citep{sora}       & 10.00 \\
Pika 1.0~\citep{pika_1.0_2023}  &  13.00 \\
SVD~\citep{svd} & 14.80 \\
Lumiere~\citep{lumiere}   & 19.00 \\
VideoPoet~\citep{videopoet} & 20.30 \\
Wan2.1~\citep{wan2025}    & 20.89 \\
Runway Gen 3~\citep{runway_gen3_alpha_2024}     & 22.80 \\
CogVideoX~\citep{cogvideox} & 26.22 \\
MAGI-1~\citep{magi1}    & 30.23 \\
\midrule
MAGI-1 + VJEPA-2 (ours) & \textbf{36.86} (+6.63) \\
\bottomrule
\end{tabularx}
\label{tab:physicsiq}
\end{table}

{
    \small
    \bibliographystyle{ieeenat_fullname}
    \bibliography{main}
}

\end{document}